train a and b are on parallel tracks. Train a is going 100 miles per hour and reaches station A at 4:15am. Train b is going 110 miles per hour and reaches station A at 4:45am. At what point will train b catch train a?
Answer by checkley77(12569)

D=RT
D=100T
D=110(T-.5)
100T=110(T-.5
100T=110T-55
100T-110T=-55
-10T=-55
T=-55/-10
T=5.5 HOURS AFTER 4:15 TRAIN B WILL CATCH TRAIN A.
PROOF:
100*5.5=110(5.5-.5)
550=110*5
550=550